# ACE-RTL: When Agentic Context Evolution Meets RTL-Specialized LLMs

### v1

@@ 3.1. Generator: An RTL-Specialized LLM
Using this large-scale RTL dataset, we apply supervised fine-tuning (SFT) (Chung et al., 2024) to obtain the RTL-specialized model. The resulting model internalizes hardware domain knowledge and RTL semantics, enabling it to generate correct RTL code across diverse design tasks.
3.2. Reflector: A Frontier Reasoning LLM
While the Generator produces RTL code grounded in domain knowledge, the Reflector analyzes simulation feedback and guides refinement. Frontier generic LLMs excel at interpreting specifications, buggy code, and error reports, making them well suited for this role. In ACE-RTL, the Reflector is built upon Claude4-Sonnet to interpret verifiable feedback and generate structured fix strategies.
@@ 3.2. Reflector: A Frontier Reasoning LLM
At each iteration, the RTL code produced by the Generator is compiled and simulated using iverilog to verify functional correctness. When simulation fails, the Reflector is triggered to analyze the feedback. Automated scripts invoke the simulator and convert the raw logs into a structured format that captures the error message and the expected versus actual signal behavior. Claude4-Sonnet then serves as the Reflector’s reasoning engine: it integrates this parsed information with the design specification and the buggy RTL code to produce a structured diagnostic report. The report provides a concise explanation of the root cause together with high-level fix guidance describing how to resolve the primary failure. This structured output is then passed to the Coordinator for context evolution. Notably, the entire process depends only on iverilog, a lightweight and open-source simulator, rather than complex proprietary toolchains. This design choice makes the system portable, easy to maintain, and readily adaptable to diverse RTL design tasks.
3.3. Coordinator: A Context Evolution Engine
@@
The Coordinator performs two key functions. First, it aggregates debugging history across iterations, tracking which errors were identified, what fixes were suggested, and whether those fixes resolved the issues. This structured memory preserves useful progress and prevents regressions, allowing the Generator to focus on remaining failure cases rather than revisiting already-corrected logic.
@@
Second, the Coordinator monitors progress across iterations. If the same error persists over multiple rounds, it recognizes that the current code trajectory may be difficult to repair and triggers a restart mechanism, discarding the current implementation and prompting the Generator to produce a fresh RTL design from the specification. This leverages LLM stochasticity to explore alternative implementations that is easier to correct.
@@
By combining structured history aggregation with an adaptive restart mechanism to construct a self-evolving context, the Coordinator keeps ACE-RTL both stable and flexible. It preserves useful debugging insights when progress is made and triggers a controlled reset when stagnation occurs, enabling more efficient progress toward functionally correct RTL code.
3.4. Parallel Scaling in ACE-RTL
Although the iterative debugging framework in ACE-RTL effectively improves code correctness, it may require a large number of iterations to converge when starting from an unfavorable initial RTL generation. To address this inefficiency, we introduce a parallel scaling strategy that exploits the inherent stochasticity of LLM-based code generation to accelerate convergence.
@@ 3.4. Parallel Scaling in ACE-RTL
Instead of relying on a single debugging trajectory, the parallel scaling strategy launches multiple ACE-RTL processes simultaneously, as shown in Figure 2(b). Each process begins with a distinct RTL implementation generated by the Generator from the same input specification. These processes then proceed separately through the iterative loop of the Generator, Reflector, and Coordinator components, performing debugging and refinement in parallel.
@@
During execution, all processes are monitored concurrently. Once any process produces an RTL implementation that passes all functional test cases, the system immediately terminates all remaining processes and outputs the correct code. If none of the processes achieve functional correctness within the predefined maximum number of iterations, all are stopped automatically.
This strategy leverages the diversity introduced by LLM stochasticity, as some initial generations naturally require far fewer debugging steps than others. By exploring multiple debugging paths in parallel, the system reduces the worst-case iterations needed to reach a correct solution. Consequently, the parallel scaling strategy largely reduces the overall iteration count, as confirmed in Figure 6.
4. Experiment
We conduct a comprehensive evaluation of ACE-RTL, comparing its performance against $14$ competitive baselines on the CVDP benchmark. In addition, we present three case studies illustrating the effectiveness of key components within ACE-RTL. Besides, we also demonstrate how the proposed parallel scaling strategy largely reduces the number of iterations required to reach correct solutions.
4.1. Experimental Setup
Baselines
We evaluate ACE-RTL against $13$ SoTA baselines, including latest large open-source models Llama4-Maverick (AI, 2024), DeepSeek-v3.1 (Liu et al., 2024a), DeepSeek-R1 (Guo et al., 2025), Kimi-K2 (Team et al., 2025), and Qwen3-Coder-480B (Yang et al., 2025). Frontier proprietary models comprise 4o-mini (Hurst et al., 2024), GPT-5 (OpenAI, 2025), and Claude4-Sonnet (Anthropic, 2025a). For RTL models, we include RTLCoder-v1.1, CodeV (Zhao et al., 2025), OriGen (Cui et al., 2024a), CraftRTL, and ScaleRTL. Besides, given that ScaleRTL extends ScaleRTL via test-time scaling and achieves performance comparable to prior RTL agentic methods such as VerilogCoder and MAGE, we select it as the representative agentic baseline for evaluation on CVDP. Our variants are: (1) ACE-RTL-Generator, our RTL-specialized model, and (2) ACE-RTL, which integrates all components of this work. (3) ACE-RTL (Claude4), which uses the same agentic framework as ACE-RTL but replaces our specialized Generator with Claude4-Sonnet.
CVDP Benchmark
We adopt the CVDP-v1.0.2 benchmark as our primary evaluation suite, which includes four representative RTL design tasks: code completion (CVDP-cid002), specification-to-RTL generation (CVDP-cid003), code modification (CVDP-cid004), and code debugging (CVDP-cid016). For the latter two tasks, the input prompt includes a basic or buggy RTL implementation that is intended to be refined according to the given specification. Together, these tasks capture the core challenges faced by RTL designers in practice, covering both generative and corrective design scenarios. The benchmark contains $94$, $78$, $55$, and $35$ problems under the four categories respectively, covering a diverse range of hardware domains including arithmetic units, on-chip communication protocols, memory hierarchies, DSP kernels, and control logic. We leave testbench generation and subjective evaluation categories in CVDP to our future work. Each problem is evaluated across five independent runs. We report two complementary metrics: (1) Pass@1, the average success rate across independent runs (Liu et al., 2023b), and (2) Agentic Pass Rate (APR), defined as the ratio of unique solved problems to the total number of problems. Since agentic methods inherently perform multiple reasoning and correction iterations, comparing them against standalone LLMs solely using Pass@1 can be misleading. Therefore, we introduce APR to better capture a model’s overall problem-solving coverage across multiple trials, providing a fairer assessment of agentic approaches.
Training and Inference
ACE-RTL-Generator is finetuned from Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) using 32 compute nodes, each equipped with 8 NVIDIA A100 GPUs. We adopt tensor, pipeline, and context parallelism to enable efficient large-scale fine-tuning. Training is performed for three epochs with a context window of 32,76832{,}768 tokens and a global batch size of $128$, using a cosine-annealing learning-rate scheduler. For inference, ACE-RTL uses five parallel processes under our parallel scaling strategy, with each process limited to 30 iterations and the Generator configured with a temperature of $1.2$. The finetuned ACE-RTL-Generator is hosted via vLLM for low-latency API inference (Kwon et al., 2023), while other components leverage Claude4-Sonnet through its official API calls.
@@ 4.1. Experimental Setup
Table 1. Functional correctness on the CVDP benchmark. Each benchmark reports percentage scores for Pass@1 and Agentic Pass Rate (APR). APR is defined as the ratio of unique solved problems to the total number of problems. For agentic methods, we omit Pass@1 scores (shown as "—") because their Pass@1 equals their APR. We use bold to denote the best results achieved by our methods and underline to mark the strongest baseline results within each CVDP task category.
@@
GPT-5 $36.17$ $\underline{39.36}$ $42.31$ $47.44$ $\underline{43.64}$ $45.45$ $\underline{54.28}$ $\underline{60.00}$
Claude4-Sonnet $\underline{37.94}$ $\underline{39.36}$ $\underline{49.49}$ $\underline{51.28}$ $42.91$ $\underline{49.09}$ $51.43$ $54.29$
@@
ScaleRTL-32B $29.79$ $35.90$ $32.73$ $40.00$
Ours ACE-RTL-Generator $\mathbf{39.57}$ $40.43$ $\mathbf{49.74}$ $52.56$ $\mathbf{65.09}$ $67.27$ $\mathbf{56.00}$ $57.14$
ACE-RTL (Claude4) $\mathbf{80.85}$ $89.74$ $81.82$ $88.57$
@@
4.2. Main Results
@@
As shown in Table 1, most prior RTL-specialized models struggle on CVDP, despite their strong performance on earlier RTL benchmarks. This suggests that models trained mainly on shorter and simpler RTL data generalize poorly to the more complex and realistic design challenges posed by CVDP. In contrast, recent frontier proprietary models such as GPT-5 and Claude4-Sonnet show notable improvements and set strong new baselines, benefiting from broader RTL exposure during large-scale pretraining and stronger reasoning capabilities on complex coding tasks.
@@
Our RTL-specialized model, ACE-RTL-Generator, further advances the RTL coding frontier, outperforming all baselines on most tasks. It achieves a Pass@1 of $65.09\%$ on Code Modification (cid004), surpassing the SoTA baseline GPT-5 by $21.45\%$. This strong performance stems from our high-quality modification data, created by extracting basic RTL code segments and training the model to augment them with missing features, as discussed in Section 3.1.
By integrating our specialized Generator with the full ACE framework, ACE-RTL achieves the best overall performance on CVDP, with up to a $44.87\%$ APR improvement over SoTA baselines. ACE-RTL also outperforms the ACE-RTL (Claude4) variant, showing that an RTL-specialized model is more effective than relying solely on a frontier generic LLM. Together, these results validate the effectiveness of ACE-RTL, which iteratively evolves contextual information to guide the specialized model toward accurate RTL coding.
4.3. Case Studies
To further demonstrate the effectiveness of our approach, we present three representative case studies. The first case showcases how our RTL-specialized model (ACE-RTL-Generator) exhibits a deeper grasp of hardware design principles and canonical RTL patterns compared to generic LLMs. The second case illustrates how ACE-RTL-Reflector elevates repeated local failures into functional reasoning, enabling it to diagnose structural errors. The third case demonstrates how ACE-RTL-Coordinator identifies unproductive progress and strategically triggers restarts, enabling the system to escape local minima and reach correct RTL implementations. Notably, the examples in the latter two case studies require the highest number of iterations to reach correct solutions in our experiments.
Refer to caption Placeholder.
Figure 3. Case Study I — RS232 Transmitter Module.
Case Study I — This case study examines the RS232 Transmitter module, which performs serial data transmission driven by precise baud-rate timing. A central component is the baud-rate generator, which produces periodic pulses that step the transmitter state machine through start, data, and stop bits. Correct behavior relies on an accurate fractional clock divider to handle the non-integer ratio between the system clock and the target baud rate.
As shown in Figure 3, Claude4-Sonnet fails to produce a functionally correct implementation due to misunderstanding a fundamental RTL design pattern. Its code couples the accumulator and pulse into a single assignment, thereby losing the overflow-based pulse signal and producing unstable timing behavior. In contrast, our RTL-specialized model (i.e., ACE-RTL-Generator) correctly implements an accumulator-based divider, updating only the lower bits and using the most significant bit as a stable pulse output. This solution reflects an understanding of canonical hardware timing structures, which our model learns from curated RTL training data.
Placeholder.
Figure 4. Case Study II.
Placeholder.
Figure 5. Case Study III.
Placeholder.
Figure 6. Effect of parallel scaling.
@@ 4.3. Case Studies
     Case Study II — This case study analyzes the 64b/66b Decoder module, which performs bit-aligned extraction of mixed-mode data streams. Figure 6 shows that ACE-RTL quickly improves functional correctness but remains stuck for many iterations before passing the final test case. ACE-RTL-Reflector identifies this stagnation and analyzes the discrepancies between expected and actual outputs. By comparing these discrepancies against the provided specification, the Reflector infers that the decoding logic requires an implicit alignment transformation that is not stated in the specification. It then formulates high-level guidance describing this missing behavioral requirement, which directs the Generator to update the extraction logic accordingly, after which the design immediately passes all remaining tests.
This case illustrates how ACE-RTL-Reflector interprets systematic failure signals to detect specification gaps and elevate repeated local adjustments into understanding the implied design intent, enabling the system to resolve design-level issues that cannot be addressed through isolated fixes.
@@
Case Study III — This case study analyzes the Clock Jitter Detection module, which evaluates clock stability by comparing measured edge intervals against a timing threshold. As shown in Figure 6, ACE-RTL makes early progress but soon enters a long plateau where all attempts fail. The stagnation occurs because each iteration repeats the same flawed behavior: performing jitter detection before a valid baseline interval is established. Although the Generator produces many syntactic variants, they all retain this early-trigger issue and therefore cannot satisfy the required timing semantics.
@@
ACE-RTL-Coordinator recognizes this stalled trajectory by observing that many consecutive iterations exhibit the same assertion pattern with no meaningful improvement. Consequently, the Coordinator triggers a restart and distills high-level insights from the failed attempts, such as the need to validate the first measurement interval before performing detection. It then instructs the Generator to regenerate the design from the specification using these clarified constraints.
@@
The restart produces a simpler and more interpretable implementation that shifts the logic toward proper interval validation. After two restarts, the Generator introduces a validity flag that delays jitter detection until the first full measurement interval has been observed, which immediately restores correctness. This case shows how ACE-RTL-Coordinator identifies prolonged stagnation, extracts useful design insights, and uses restart mechanism to move the system toward correct RTL implementations.
4.4. Efficacy of Parallel Scaling
Figure 6 illustrates the impact of the proposed parallel scaling strategy on reducing the iteration count required for convergence across multiple CVDP categories. The results clearly show that launching multiple independent ACE-RTL instances in parallel significantly accelerates the debugging process without compromising correctness. Specifically, the average iteration count drops from $11.33$ to $3.95$ for cid002, from $11.25$ to $4.23$ for cid003, from $9.25$ to $3.75$ for cid004, and from $13.36$ to $4.37$ for cid016, yielding speedups of $2.87\times$, $2.66\times$, $2.47\times$, and $3.06\times$, respectively.
This improvement demonstrates the efficacy of the parallel scaling design: by exploring diverse debugging trajectories derived from different initial generations, ACE-RTL avoids being constrained by any single unfavorable starting point. Instead, at least one parallel process typically converges rapidly, allowing other trajectories to be terminated early. Consequently, the system achieves nearly a $3\times$ runtime reduction to explore a correct solution.
4.5. Results on VerilogEval

### v2

@@ III-A Generator: An RTL-Specialized LLM
Using this large-scale RTL dataset, we apply supervised fine-tuning (SFT) [6] to obtain the RTL-specialized model. The resulting model captures rich hardware domain knowledge and RTL semantics, enabling accurate code generation across diverse design tasks. Moreover, training on editing and debugging supervision enables the model to effectively analyze, refine, and correct RTL implementations, making it well suited for deployment within our agentic framework.
III-B Reflector: A Frontier Reasoning LLM
While our Generator produces and refines implementations within the agentic loop, effective iteration also requires interpreting simulation outputs and identifying corrective actions. We therefore introduce the Reflector, a generic frontier LLM (Claude4-Sonnet) that analyzes simulation feedback and long-context error reports to guide subsequent refinement. Processing such complex information benefits from the broader reasoning capability and context capacity of frontier models.
@@ III-B Reflector: A Frontier Reasoning LLM
At each iteration, we invoke iverilog once to compile and execute the candidate RTL against the benchmark test harness. When the run fails, the resulting logs are parsed into a structured report capturing the error message together with the expected-versus-actual signal behavior. Claude4-Sonnet then serves as the Reflector’s reasoning engine: it combines this structured feedback with the design specification and the current RTL implementation to produce a concise diagnosis and high-level fix guidance. This output is then passed to the Coordinator to construct the evolving context for subsequent refinement steps. Notably, the entire process depends only on iverilog, a lightweight and open-source simulator, rather than complex proprietary toolchains.
III-C Coordinator: A Context Evolution Engine
@@
The Coordinator performs two key functions. First, it aggregates debugging history across iterations, tracking identified errors, suggested fixes, and their outcomes. This history is incorporated into the evolving context, preserving useful progress and preventing regressions so that the Generator can focus on unresolved failures.
@@
Second, the Coordinator monitors progress across iterations. It uses Claude4-Sonnet to examine the debugging trajectory and determine whether the current search has stalled. When the same primary failure persists across several iterations without meaningful improvement, the Coordinator triggers a restart mechanism, discarding the current implementation and prompting the Generator to produce a fresh RTL design. This leverages LLM stochasticity to explore alternative implementations that are easier to correct.
@@
By combining cross-iteration history aggregation with an adaptive restart mechanism, the Coordinator keeps the evolving context informative and flexible. It preserves useful debugging insights when progress is made and triggers controlled resets during stagnation, enabling efficient convergence to correct RTL code.
Notably, from the Generator’s perspective, the self-evolving context resembles an augmented specification that describes failure symptoms or required modifications. Because the Generator is trained on both editing and debugging tasks, it is naturally capable of interpreting such context, without requiring additional model adaptation.
III-D Parallel Scaling in ACE-RTL
@@
Although the iterative debugging framework in ACE-RTL effectively improves code correctness, it may require a large number of iterations to converge when starting from an unfavorable initial RTL generation. To address this inefficiency, we introduce a parallel scaling strategy that exploits the inherent stochasticity of LLM-based code generation to accelerate convergence. Instead of relying on a single debugging trajectory, the parallel scaling strategy launches multiple ACE-RTL processes simultaneously, as shown in Figure 2(b). Each process begins with a distinct RTL implementation generated by the Generator from the same input specification. These processes then proceed separately through the iterative loop of the Generator, Reflector, and Coordinator components, performing debugging and refinement in parallel.
@@
During execution, all processes are monitored concurrently. Once any process produces an RTL implementation that passes all functional test cases, the system immediately terminates all remaining processes and outputs the correct code. If none of the processes achieve functional correctness within the predefined maximum number of iterations, all are stopped automatically. This strategy leverages LLM generation diversity to reduce time-to-first-success with a trade-off in total inference-time compute. In Section IV-E, we evaluate parallel scaling from both convergence and inference-cost (i.e., token usage) perspectives. Our goal is not to claim lower total compute, but to show that parallel scaling provides a favorable speed-cost trade-off.
IV Experiment
IV-A Experimental Setup
@@
Table 1: Functional correctness on CVDP. We report Pass@1 and Agentic Pass Rate (APR). For agentic methods, Pass@1 is omitted (“—”) since it equals APR. Bold indicates our best results, and underline the strongest baseline in each category.
@@
GPT-5 $36.17$ $39.36$ $42.31$ $47.44$ $\underline{43.64}$ $45.45$ $\underline{54.28}$ $60.00$
Claude4-Sonnet $\underline{37.94}$ $39.36$ $\underline{49.49}$ $51.28$ $42.91$ $49.09$ $51.43$ $54.29$
@@
RTL Agents MAGE (Claude4) $44.68$ $53.85$ $50.91$ $60.00$
MAGE (ACE-RTL-Generator) $\underline{46.81}$ $\underline{55.13}$ $\underline{70.91}$ $\underline{62.86}$
Ours Generator (standalone) $\mathbf{39.57}$ $40.43$ $\mathbf{49.74}$ $52.56$ $\mathbf{65.09}$ $67.27$ $\mathbf{56.00}$ $57.14$
@@
Baselines
We evaluate ACE-RTL against $14$ SoTA baselines, including frontier open-source models Llama4-Maverick [1], DeepSeek-v3.1 [18], DeepSeek-R1 [12], Kimi-K2 [29], and Qwen3-Coder-480B [32]. Frontier proprietary models comprise 4o-mini [15], GPT-5 [26], and Claude4-Sonnet [4]. For RTL models, we include RTLCoder-v1.1, CodeV [36], OriGen [7], CraftRTL, and ScaleRTL. In addition, as VerilogCoder relies on a custom Verilog parsing tool that is not readily portable to SystemVerilog designs in CVDP, we adopt MAGE as the RTL agentic baseline. To ensure a fair comparison, we use Claude4-Sonnet as the underlying LLM in MAGE, which is the same generic model used in ACE-RTL. Our variants are: (1) Generator, our RTL-specialized model; (2) ACE-RTL, the full system integrating the specialized Generator with the Reflector and Coordinator.
CVDP Benchmark
We adopt the CVDP-v1.0.2 benchmark as our primary evaluation suite, which includes four representative RTL design tasks: code completion (CVDP-cid002), specification-to-RTL generation (CVDP-cid003), code modification (CVDP-cid004), and code debugging (CVDP-cid016). Together, these tasks capture the core challenges faced by RTL designers in practice, covering both generative and corrective design scenarios. The benchmark contains $94$, $78$, $55$, and $35$ problems under the four categories respectively, covering a diverse range of hardware domains including arithmetic units, on-chip communication protocols, memory hierarchies, DSP kernels, and control logic. We leave testbench generation and subjective evaluation categories in CVDP to our future work. Each problem is evaluated across five independent runs. We report two complementary metrics: (1) Pass@1, the average success rate across independent runs [20], and (2) Agentic Pass Rate (APR), defined as the ratio of unique solved problems to the total number of problems. Since agentic methods inherently perform multiple reasoning and correction iterations, comparing them against standalone LLMs solely using Pass@1 can be misleading. Therefore, we introduce APR to better capture a model’s overall problem-solving coverage across multiple trials, providing a fairer assessment of agentic approaches.
Training and Inference
ACE-RTL-Generator is finetuned from Qwen2.5-Coder-32B-Instruct [14] using 32 nodes with 8 NVIDIA A100 GPUs each. We employ tensor, pipeline, and context parallelism for efficient large-scale training, and adopt sequence packing over the training data to improve token utilization and speed up the training process. The model is trained for three epochs with a 32,768-token context window and a global batch size of 128 using a cosine-annealing learning rate. The total training process takes approximately 10K GPU-hours, which is a relatively modest cost for training a strong domain-specialized model, especially compared with frontier LLM training, which typically requires multiple orders of magnitude more compute. During inference, ACE-RTL runs five parallel processes under our parallel scaling strategy, each limited to 30 iterations with a temperature of 1.2. The finetuned Generator is served via vLLM [17], while other components use Claude4-Sonnet through its official API.
Figure 3: Comparison of frontier generic and RTL-specialized LLMs.
IV-B Main Results
@@
Table 1 shows prior RTL LLMs struggle on CVDP despite performing well on earlier RTL benchmarks, suggesting limited generalization from shorter and simpler RTL data to the more complex designs in CVDP. In contrast, frontier models such as GPT-5 and Claude4-Sonnet show clear improvements and establish strong new baselines, benefiting from broader RTL exposure during large-scale pretraining and stronger reasoning on complex coding tasks.
@@
Our RTL-specialized model, ACE-RTL-Generator, further advances the RTL coding frontier, outperforming all standalone model baselines on most tasks. It achieves a Pass@1 of $65.09\%$ on Code Modification (cid004), surpassing the SoTA baseline GPT-5 by $21.45\%$. This strong performance stems from our high-quality modification data, created by extracting basic RTL code segments and training the model to augment them with missing features, as discussed in Section III-A.
Figure 3 compares standalone models in terms of average CVDP Pass@1 across four categories versus model size. In addition to achieving the highest standalone pass rate, ACE-RTL-Generator also lies beyond the Pareto frontier formed by prior baselines, indicating a particularly favorable trade-off between model scale, which is broadly correlated with training cost, and functional correctness.
For agentic baselines, we evaluate two MAGE variants, using either Claude4 or ACE-RTL-Generator as the underlying RTL generation model. In both cases, MAGE achieves only marginal gains over the corresponding standalone model. We attribute this mainly to the absence of a self-evolving context that preserves useful learnings from prior iterations. Without such accumulated debugging history, the agentic loop has limited ability to build on past failures, making it harder to identify root causes and apply targeted fixes over time. A secondary limitation is that MAGE depends on a testbench agent whose outputs can be inaccurate, even when given the golden CVDP test harness, because translating cocotb-based CVDP harnesses into SystemVerilog adds another layer of difficulty on top of the well-known challenge of testbench generation for complex designs [28]. These results suggest that prior SoTA agentic baselines still struggle to generalize beyond earlier, simpler benchmarks.
Overall, ACE-RTL achieves the best performance on CVDP, with up to a $41.02\%$ APR improvement (on cid003) over all baselines. These results show that combining our RTL-specialized model with iterative context evolution is highly effective for complex RTL tasks.
IV-C Ablation Study
Figure 4 reports the averaged APR of several ablated variants of ACE-RTL across the four CVDP categories. Relative to the standalone Generator, ACE-RTL w/o Coordinator already yields a clear improvement, confirming the value of simulation-guided iterative refinement. The largest gain, however, comes from enabling the Coordinator in ACE-RTL w/o restart, which maintains an evolving context across iterations using debugging history from prior attempts. This result shows that preserving cross-iteration history is a key ingredient in our agentic system, allowing the Generator to build on prior progress rather than repeatedly revisiting similar failures.
Adding the restart mechanism brings another clear improvement over ACE-RTL w/o restart, indicating that preserving the evolving context alone is not always sufficient. Some failed trajectories remain difficult to recover through incremental local fixes and instead benefit from regenerating a fresh implementation under the accumulated debugging context. This behavior is also illustrated in Figure 7, where restart helps the system escape prolonged stagnation. Finally, ACE-Claude4 uses the same ACE framework but replaces our RTL Generator with Claude4. Although it achieves strong performance, the full ACE-RTL still performs best overall, demonstrating that agentic context evolution and RTL specialization are complementary.
Figure 4: Ablation study on ACE-RTL components.
Figure 5: Case Study I.
Figure 6: Case Study II.
Figure 7: Effect of parallel scaling.
IV-D Case Studies
To further demonstrate the effectiveness of ACE-RTL, we present two representative case studies. The first illustrates how ACE-RTL moves beyond repeated local fixes to uncover the underlying functional mistake in RTL. The second illustrates how ACE-RTL identifies unproductive search trajectories and strategically restarts, enabling it to escape local minima and reach a correct RTL implementation.
@@ IV-D Case Studies
     Case Study I — This case study analyzes the 64b/66b Decoder module, which performs bit-aligned extraction of mixed-mode data streams. Figure 7 shows that ACE-RTL quickly improves functional correctness but remains stuck for many iterations before passing the final test case. Our Reflector identifies this stagnation and analyzes the discrepancies between expected and actual outputs. By comparing these discrepancies against the provided specification, the Reflector infers that the decoding logic requires an implicit alignment transformation that is not clearly stated in the specification. It then formulates high-level guidance describing this behavioral requirement, which directs the Generator to update the extraction logic accordingly, after which the design immediately passes all remaining tests.
@@
Case Study II — This case study analyzes the Clock Jitter Detection module, which evaluates clock stability by comparing measured edge intervals against a timing threshold. As shown in Figure 7, ACE-RTL makes early progress but soon enters a long plateau where all attempts fail. The stagnation occurs because each iteration repeats the same flawed behavior: performing jitter detection before a valid baseline interval is established. Although the Generator produces many syntactic variants, they all retain this early-trigger issue and therefore cannot satisfy the required timing semantics.
@@
Our Coordinator recognizes this stalled trajectory by observing that many consecutive iterations exhibit the same assertion pattern with no meaningful improvement. Consequently, the Coordinator triggers a restart and distills high-level insights from the failed attempts, such as the need to validate the first measurement interval before performing detection. It then instructs the Generator to regenerate the design from the specification using these clarified constraints.
@@
The restart yields a simpler implementation that properly validates the measurement interval. After two restarts, the Generator introduces a validity flag that delays jitter detection until the first full interval is observed, restoring correctness. This example illustrates how the Coordinator detects stagnation and uses a restart mechanism to guide the system toward a correct RTL implementation.
Table 2: Impact of PS (parallel scaling) on convergence speed and inference cost, averaged over the four CVDP categories.
Setting APR $\uparrow$ Iter. $\downarrow$ Speedup $\uparrow$ Rel. Token $\downarrow$
w/o PS 89.84 11.30 $\times 1.00$$\times$
[…]
IV-E Efficacy of Parallel Scaling
[…]
IV-F Results on VerilogEval